Meeting Notes for 17 Feb 2010

Present: Phong, Mihael, Liz, TomJ, TomL
Absent: Marge, Bob

Action Items

  • Make a CMS roadmap


  • The planned data upgrade on Monday went very badly; the IBM storage server refused to have anything to do with our new drives, our proposed backup server was confused by the new drives and, worse, the IBM server's drive backplane appeared to die. frown, sad smile Trouble ticket opened with Big Blue.
  • New network switch is working
  • Need global news item (not per-lab) for downtime and such.


  • Blessing meeting tomorrow at ~1000. See attachment for design proposal.
  • UI Design still need to be hashed out?
  • Background information for users need to be written up


  • In progress
  • Showing interesting data anomalies (bias changes, null data)
  • LIGO eLab workshop for homeschool students on 6, 20 March.
    • Bluestone on PROD gets confused with the new data; should put a bounds on the date filter.
    • Alert Dale on this!


  • Putting a masterclass together for ~200 people.
  • Education side has made a lot of progress
    • Get the LOs to Jean - feedback should be in by next week's teleconference.
  • Get back to Dave Barney, et. al. on our proposal

-- Main.PhongNguyen - 17 Feb 2010
Topic attachments
I Attachment Action Size Date Who Comment
Fwd__Data_Blessing_Design_II.txttxt Fwd__Data_Blessing_Design_II.txt manage 6 K 2010-02-17 - 17:13 Main.UnknownUser Data Blessing Design Document
Topic revision: r21 - 2010-02-17, PhongNguyen
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback